OpenStudy (loser66):

Check my stuff, please. z = i w = -1 +i zw = -1-i

OpenStudy (anonymous):

Ill help Brb

OpenStudy (loser66):

arg z = pi/2 arg w = 3pi/4 arg (zw) = 5pi/4, what is wrong with them?

OpenStudy (anonymous):

Nevermind I cant help But I give Medal And Fan Sry

OpenStudy (loser66):

@Empty

OpenStudy (empty):

Nothing is wrong with them, arg(z)+arg(w)=arg(zw) in general.

OpenStudy (loser66):

what is log (zw) ?

OpenStudy (loser66):

I have to prove log (zw) \(\neq \) log z + log w because of their argument from both sides are not the same. But I don't see it :(

OpenStudy (empty):

Depends, it's multivalued since \[e^{i \theta} = e^{i(\theta + 2 \pi)}\] we can do this an integer amount of times + or - so we have: Every complex number has a polar form: \[z=re^{i(\theta + 2 \pi n)}\] \[\log(z) =\log (re^{i(\theta + 2 \pi n)}) = \log (r) + i(\theta + 2 \pi n)\]

OpenStudy (anonymous):

your'e correct

OpenStudy (empty):

http://math.stackexchange.com/a/927129/207119

OpenStudy (empty):

You're trying to prove a false statement as far as I can tell.

OpenStudy (loser66):

(zw) = -1-i, hence its argument is 5pi/4

OpenStudy (loser66):

and this argument is = arg z + arg w, right?

OpenStudy (empty):

\[\frac{\pi}{2}=\frac{2 \pi}{4}\] \[\frac{2 \pi}{4} + \frac{3 \pi}{4} = \frac{5 \pi}{4} \]

OpenStudy (loser66):

So, for those numbers, z and w. The statement is true. right? log (zw) = log z + log w

OpenStudy (loser66):

Hence, to prove the statement is wrong, I have to pick other z, w, right?